Given that 𝐢 = 100 + 0.8π‘Œ; 𝑀 = 150 + 0.2π‘Œ; 𝐼 = 100 π‘Žπ‘›π‘‘ 𝑋 = 350,


a) Determine the equilibrium income


b) Show the results on two graphs with (i) Injections and leakages on the vertical axis


(ii) Net injections and net domestic leakages on the vertical axis


c) Determine the new equilibrium and show on graph when there is;


i. An increase in X and I of 200


ii. A decrease in M of 100


iii. A decrease in S and M of 100



1
Expert's answer
2022-02-28T11:34:20-0500

A)Most simply, the formula for the equilibrium level of income is when aggregate supply (AS) is equal to aggregate demand (AD), where AS = AD.


Adding a little complexity, the formula becomes Y = C + I + G(X-M), where Y is aggregate income, C is consumption, I is investment expenditure, and G is government expenditure

"Y=100+0.8Y+100+(350-150)"

"Y-0.8Y=400"

"Y(0.2)=400"

"Y=2000"

B)


ii)


c)

i)

Y=3000

ii)

Y=2500

iii)

Y=2000